Supplemental Operations Report – December 2nd

MDRS Supplemental Operations Report 2-Dec-2019

Name of person filing report: Shannon Rupert

Reason for Report: Routine update

Non-nominal systems: RAM electrical

Action taken for non-nominal systems: Troubleshooting, no solutions. We now have our second incidence of electrical work done without permission and with no reportage. Someone has rewired the east side electric in the RAM that had been incorrectly wired when it was built. In the meantime, this appears to have damaged the wiring to the rover chargers and now they are not functional, and continually trip when the rovers are plugged in. This appears to be an issue separate from the initial problem of incorrectly installed GFI sockets and insufficient breaker strength. We intend to take up this issue in the next few weeks.

People doing things without notifying us is a common problem. Yesterday the power system had been adjusted, but not reset, so I was able to easily readjust it back where is needed to be. Later it was discovered that the crew did not have water because someone in the crew had turned it all off. But no one will say that they did it. It is the same reluctance to speak up that happened when the heat in the GreenHab was turned off, it was physically impossible for it to have happened without someone doing it, but no one wanted to say that they had done it. Until we get to the point where everyone is willing to practice full disclosure, we cannot and will not have a safe and successful Mars mission. This is something we need to get used to doing here.

Generator check, note if oil and coolant added: coolant added. Suspect
oil leak appears to not be a problem.

ScienceDome AC unit, on and operational: nothing to report

Solar— Running nominally with nighttime generator backup

Diesel Reading – % not checked

Propane Reading – not checked

Ethanol Free Gasoline – n/a gallons. Now stored in the shed.

Water (director’s tank) – n/a gallons

Water (static tank) – n/a gallons

Water (GreenHab) – n/a gallons

Water (loft) – nothing to report

Static to Loft Pump used – yes

Water Meter: n/a

Toilet tank emptied: nothing to report

Deimos rover used: Yes, left today for repairs in Richfield


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: Yes, operation nominal


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Yes, operation nominal


Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: Yes, left today for brake repairs in Richfield


Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: Yes, operation nominal.


Beginning charge:

Ending charge:

Currently charging:

Notes and action taken on rovers: Rovers were cleaned of mud by the crew.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3). No. 300 still in town.

Reason for use: n/a

Oil Added? n/a

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used: none

Notes and action taken on ATVs: n/a

HabCar running? If used, why, where? Yes, used to town for water
runs and supplies

CrewCar running, If used why, where? Yes, currently onsite and used to transport crews. Noted that the check engine light and service tire monitor light were not serviced when the vehicle was in the shop. Trying to determine if we paid for that or not, will probably return to dealer in Grand Junction if not.

Campus wide inspection, if action taken, what and why? Burn barrel was finally filled with holes and will be put into service back in the Outpost so we won’t have to be in sight of the crew. Old burn site will be restored to pre-burn status as soon as possible.

Evidence of rodents, where and action taken: We appear to have one brave little beast living in Little Outpost (formerly the Intern Trailer). So far he has just mocked Atila’s attempts at capture and keeps eating the cookies.

Summary of internet: All nominal.

Summary of suits and radios: Changes to radio channels has made a great improvement in COMMS.

Summary of Hab operations: New water heater has been installed. This is not an on demand model. Those have proven to be unsuited to our water system. This is a 6 gallon RV tank model, installed on the lower deck above the rear airlock, where a water heater had been
installed before. It is both gas and electric. Testing it in electric mode tonight and tomorrow we will install the 12 V power and switch for propane.

Propane was reconnected to the stove and hot plate was removed.

Grey vents from former hot water heater were reinstalled on the exterior of the Hab, covering the two holes.

Water tank heater installed and operational.

Summary of Outpost operations: Emptying the blackwater tank of the Little Outpost the first time was only marginally successful. We need to install a permanent line. Shower drain repaired.

Septic needs to be pumped. Riser was affixed to the tank so yard could be leveled and hole filled in. It will take a while to move all the dirt back into place but we have started on it.

Tank heater operational.

Summary of GreenHab operations: 40% shade cloth covering 50 % of Greenhab south side. A 30% shade cloth was ordered to cover the other 50%.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Electrical work required. See non-functional systems report above

Summary of any observatory issues: Atila and Peter are troubleshooting a cable issue on the robotic observatory. Testing will determine if it has been resolved.

Summary of health and safety issues: We do not have a functional blood pressure monitor so I ordered one to arrive this week.

List and summary of projects:

Tennis balls to protect the tarps on the tunnels: Let’s just say we hate cutting tennis balls. We vow to finish soon.

Questions, concerns, supplies needed and requests: Nothing at this time.

Copyright © The Mars Society. All rights reserved. | Main Site